OpenStudy (jazzyfa30):

Help Please Write the equation y = 2x +3 in standard form.

OpenStudy (amistre64):

what does "standard" form look like?

OpenStudy (jazzyfa30):

y=mx+b

OpenStudy (amistre64):

no, that is slope intercept form :) standard form is: Ax + By = C

OpenStudy (amistre64):

so just get the x and y parts on the same side

OpenStudy (jazzyfa30):

Oh ok

OpenStudy (amistre64):

it might be worth noting that A should be positive so you might have to adjust for that

OpenStudy (jazzyfa30):

x+y=3 /2

OpenStudy (amistre64):

y = 2x +3 -2x -2x ----------- -2x+y = 3 multiply both sides by -1 to adjust for a postive x part 2x - y = -3

OpenStudy (jazzyfa30):

Oh I get it now

OpenStudy (amistre64):

it is rather simple :) what i found the most difficultest was trying to remember the names of all the line equation formats .....

OpenStudy (jazzyfa30):

That's where I get mixed up at

OpenStudy (amistre64):

good luck :)

OpenStudy (jazzyfa30):

Thanks I needed that LOL ;)